# Fanxiao Li 1, Jiaying Wu2, Tingchao Fu1, Yunyun Dong1, Bingbing Song1, Wei Zhou 1

### v1

@@ A.1 Dataset Creation Details
In section 3.2, we discussed the source data collection process of DriftBench. We outlines the key implementation details and the prompt templates used to generate the diversified and fabricated content in the DriftBench, including both text and image transformations.
@@
: Given a human written news text (OT), we prompt GPT-4o (Hurst et al. 2024) to rephrase the original text while preserving its semantics.
@@
: Given a human written news text (OT), we prompt GPT-4o (Hurst et al. 2024) to generate a false narrative by altering the original news caption.
@@
: Given an authentic image (OI), we utilize FLUX.1 Redux [dev] (Labs et al. 2025), an open-source image-to-image variant generation model, to generate diversified image. We set the guidance_scale as 5.5 and num_inference_step as 50.
@@
: Given an false narrative (FI), we utilize FLUX.1 [dev] (Labs 2024), an open-source text-to-image generation model, to generate fabricated image. We set the guidance_scale as 3.5 and num_inference_step as 50.
@@ A.2 Human Evaluation
We illustrate the annotation process for human evaluation discussed in Section 3.3 in Figure 7.
@@ A.3 Malicious Evidence Contamination
To implement malicious evidence contamination through guided retrieval, we prompt GPT-4o (Hurst et al. 2024) to generate evidence with malicious intent, such as evidence supporting or refuting news instances, and directly insert it into the original evidence list. The original evidence list is curated by crawling data from the web and selecting the top 5 pieces of evidence. To evaluate the impact of evidence contamination, we insert the generated malicious evidence into the list and use it for model inference.
@@ A.4 Error Attribution
We listed detailed error attribution taxonomy in Figure 8 discussed in Section 3.4.
@@
Model misperception drift errors: Text–Image inconsistency: The image and text do not align, leading to contradictions, such as the wrong person, place, action, or object being depicted. Text–Image ambiguous or weak match: The image is too vague or the caption is too generic, making it difficult to establish a clear and meaningful connection between them. Image and caption match, but the claim itself is fabricated: The image and caption are consistent, but the event or claim described is fabricated or never occurred (e.g., a historical event that never happened). Evidence drift errors: No supporting evidence found: The retrieval process fails to find any relevant evidence or retrieves evidence from irrelevant domains, leading to a lack of support for the claim. Misleading retrieved evidence: The retrieved evidence may have similar names or topics to the claim but is ultimately a poor or incorrect match. Mismatch in time or location between the evidence and the caption: The time or location of the retrieved evidence does not align with that presented in the caption, leading to inconsistencies. Over-reliance on low-quality or noisy evidence: The model excessively depends on low-quality or noisy evidence, which undermines the overall reliability and validity of the claim’s verification.
@@ A.5 Explanation Evaluation
To analyze the reasoning behavior of the model, we evaluated the explanations generated by different methods discussed in 3.4. The detailed evaluation dimensions and criteria is shown in Figure 9, and the results for all models are shown in the Figure 10. In this part, we also use GPT-4o (Hurst et al. 2024) to conduct our experiment.
@@ A.6 Detailed Results
We present the detailed performance of Real vs. Fake instances across all diversified types discussed in Section 4.2 in Table 4. Additionally, we report the results of robustness to evidence contamination in Table 5, as discussed in Section 4.3.

### v2

@@ A.1 Dataset Creation Details
In section Dataset Creation, we discussed the source data collection process of DriftBench. We outlines the key implementation details and the prompt templates used to generate the diversified and fabricated content in the DriftBench, including both text and image transformations.
@@
: Given a human written news text (OT), we prompt GPT-4o (DBLP:journals/corr/gpt-4o) to rephrase the original text while preserving its semantics.
@@
: Given a human written news text (OT), we prompt GPT-4o (DBLP:journals/corr/gpt-4o) to generate a false narrative by altering the original news caption.
@@
: Given an authentic image (OI), we utilize FLUX.1 Redux [dev] (labs2025flux1kontextflowmatching), an open-source image-to-image variant generation model, to generate diversified image. We set the guidance_scale as 5.5 and num_inference_step as 50.
@@
: Given an false narrative (FI), we utilize FLUX.1 [dev] (flux2024), an open-source text-to-image generation model, to generate fabricated image. We set the guidance_scale as 3.5 and num_inference_step as 50.
@@ A.2 Human Evaluation
We illustrate the annotation process for human evaluation discussed in Section Data Quality Assessment in Figure 7.
@@ A.3 Malicious Evidence Contamination
To implement malicious evidence contamination through guided retrieval, we prompt GPT-4o (DBLP:journals/corr/gpt-4o) to generate evidence with malicious intent, such as evidence supporting or refuting news instances, and directly insert it into the original evidence list. The original evidence list is curated by crawling data from the web and selecting the top 5 pieces of evidence. To evaluate the impact of evidence contamination, we insert the generated malicious evidence into the list and use it for model inference.
@@ A.4 Error Attribution
We listed detailed error attribution taxonomy in Figure 8 discussed in Section Evaluation Tasks and Metrics.
@@
Model misperception drift errors: Text–Image inconsistency: The image and text do not align, leading to contradictions, such as the wrong person, place, action, or object being depicted. Text–Image ambiguous or weak match: The image is too vague or the caption is too generic, making it difficult to establish a clear and meaningful connection between them. Image and caption match, but the claim itself is fabricated: The image and caption are consistent, but the event or claim described is fabricated or never occurred (e.g., a historical event that never happened). Model misperception drift errors: No supporting evidence found: The retrieval process fails to find any relevant evidence or retrieves evidence from irrelevant domains, leading to a lack of support for the claim. Misleading retrieved evidence: The retrieved evidence may have similar names or topics to the claim but is ultimately a poor or incorrect match. Mismatch in time or location between the evidence and the caption: The time or location of the retrieved evidence does not align with that presented in the caption, leading to inconsistencies. Over-reliance on low-quality or noisy evidence: The model excessively depends on low-quality or noisy evidence, which undermines the overall reliability and validity of the claim’s verification.
@@ A.5 Explanation Evaluation
To analyze the reasoning behavior of the model, we evaluated the explanations generated by different methods discussed in Evaluation Tasks and Metrics. The detailed evaluation dimensions and criteria is shown in Figure 9, and the results for all models are shown in the Figure 10. In this part, we also use GPT-4o (DBLP:journals/corr/gpt-4o) to conduct our experiment.
@@ A.6 Detailed Results
We present the detailed performance of Real vs. Fake instances across all diversified types discussed in Section Performance under GenAI-Driven Diversity in Table 4. Additionally, we report the results of robustness to evidence contamination in Table 5, as discussed in Section Robustness to Evidence Contamination.